2x(y+z^2)p+(2y+z^2)q=z^3
1
Expert's answer
2020-12-10T14:00:33-0500

"\displaystyle
2x(y+z^2)p+(2y+z^2)q=z^3\\

\textsf{The lagrange auxiliary equation is}\\

\frac{\mathrm{d}x}{2x(y + z^2)} = \frac{\mathrm{d}y}{2y + z^2} = \frac{\mathrm{d}z}{z^3}\\

\textsf{Choosing multipliers}\,\,\left(\frac{1}{x}, -1, -\frac{1}{z}\right)\\

\int\frac{\mathrm{d}x}{x} - \int\mathrm{d}y - \int\frac{\mathrm{d}z}{z} = 0\\

\int\frac{\mathrm{d}x}{x} - \int\mathrm{d}y = \int\frac{\mathrm{d}z}{z} \\

\ln{x} - y - C = \ln{z}\\

\ln\left(\frac{x}{z}\right) = y + C\\

\frac{x}{z} = Ae^{y}, x = Aze^{y}\\

\textsf{Comparing the last two equations}\\

\frac{\mathrm{d}y}{2y + z^2} = \frac{\mathrm{d}z}{z^3}\\

\frac{\mathrm{d}y}{\mathrm{d}z} =\frac{2y + z^2}{z^3}\\

\frac{\mathrm{d}y}{\mathrm{d}z} =\frac{2y}{z^3} + \frac{1}{z}\\

\frac{\mathrm{d}y}{\mathrm{d}z} - \frac{2y}{z^3} = \frac{1}{z}\\

\textsf{Integration Factor}\,(IF) = e^{-\int \frac{2}{z^3}\, \mathrm{d}z} = e^{\frac{1}{z^2}}\\


y\cdot IF = \int \frac{IF}{z}\,\mathrm{d}z\\

ye^{\frac{1}{z^2}} = \int \frac{e^{\frac{1}{z^2}}}{z} \,\mathrm{d}z\\

\textsf{Let}\,\, u = \frac{1}{z^2}, -\frac{z^3}{2}\mathrm{d}u = \mathrm{d}z\\

ye^{\frac{1}{z^2}} = \int \frac{e^{u}}{z} \cdot -\frac{z^3}{2}\mathrm{d}u \\

ye^{\frac{1}{z^2}} = \int e^u\cdot -\frac{z^2}{2}\mathrm{d}u \\

ye^{\frac{1}{z^2}} = \frac{-1}{2}\int \frac{e^u}{u}\,\mathrm{d}u \\

\textsf{The integral on the right side}\\
\textsf{can be written in terms}\\
\textsf{of the the exponential integral}\,\, \mathrm{Ei}(u)\\

ye^{\frac{1}{z^2}} = \frac{-1}{2}\int_0^u \frac{e^t}{t}\,\mathrm{d}t \\

ye^{\frac{1}{z^2}} = \frac{-\mathrm{Ei}(u)}{2}\,\, y = \frac{-e^{\frac{-1}{z^2}}\mathrm{Ei}(u)}{2}\\

y = \frac{-e^{\frac{-1}{z^2}}\mathrm{Ei}\left(\frac{1}{z^2}\right)}{2}\\

\textsf{Therefore, the solution to PDE is}\,\,\\

x = Aze^{y}, y = \frac{-e^{\frac{-1}{z^2}}\mathrm{Ei}\left(\frac{1}{z^2}\right)}{2}\\"